OpenStudy (anonymous):

When solving x^2 – 16x = 10 by completing the square what value is added to both sides of the equation?

OpenStudy (anonymous):

(x-16)x=0 (x-8)^2-74=0 x^2-16x-10=0

OpenStudy (anonymous):

64, half of the coefficient of x, squared

OpenStudy (anonymous):

in this situation it would be 8^2.. bec u take the 16 and half it, then square it....